# Introducing FOReCAst: The Future Outcome Reasoning and Confidence Assessment Benchmark

### v3

@@ Abstract
Forecasting is an important task in many domains, such as technology and economics. However existing forecasting benchmarks largely lack comprehensive confidence assessment, focus on limited question types, and often consist of artificial questions that do not align with real-world human forecasting needs. To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), [^1] Publicly availble at https://huggingface.co/datasets/MoyYuan/FOReCAst a benchmark that evaluates models’ ability to make predictions and their confidence in them. FOReCAst spans diverse forecasting scenarios involving Boolean questions, timeframe prediction, and quantity estimation, enabling a comprehensive evaluation of both prediction accuracy and confidence calibration for real-world applications.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks. Alongside these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023). Fact-based benchmarks, such as TruthfulQA (Lin et al., 2022), evaluate LLMs based on factual correctness, focusing on tasks like retrieving and verifying facts that are already known.
@@
Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy. Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information. This makes forecasting particularly difficult, as models must infer trends, assess probabilities, and adapt to new information. Several datasets have been introduced to evaluate LLMs’ forecasting capabilities. ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment. AutoCast (Zou et al., 2022) incorporates confidence intervals; however, its confidence estimates are not specifically designed for forecasting tasks. Other datasets, such as ExpTime (Yuan et al., 2024a), are artificially generated from structured data and focus on explainable event forecasting based on temporal knowledge graphs.
@@
All of these aforementioned benchmarks overlook a crucial aspect of forecasting: confidence evaluation. Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive correctness at the time of evaluation. Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events. Moreover, miscalibrated confidence can lead to poor decision-making: overconfident yet incorrect forecasts may result in costly errors, while underconfident but accurate predictions can erode trust in the model. Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
@@
To address these gaps, we present FOReCAst: Future Outcome Reasoning and Confidence Assessment. FOReCAst focuses on three distinct types of forecasting questions, shown in Table 1: (1) Boolean Questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in each of the following years?" We conduct experiments using a range of models differing in size, training objectives, and training cutoff times, and explore multiple methods for estimating model confidence. Our results reveal that forecasting remains highly challenging for current LLMs, particularly in confidence evaluation, with no consistent correlation between prediction accuracy and calibration. While larger models sometimes improve performance, the effect is inconsistent and highly task-dependent.
@@
Boolean Question Will a Frontier AI lab be established in China by 2026? Yes 0.73
@@
Table 1: Examples of forecasting questions with their resolutions and confidence scores.
@@ 2 FOReCAst: Problem Formulation
System responses in FOReCAst consist of (1) a prediction answering a question given the available information and (2) a confidence score in the prediction. This ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
Questions belong to three types. (1) Boolean Questions, which ask yes/no questions about the occurrence of future events (sometimes within a certain timeframe). Boolean questions are simple to evaluate, and they can still be surprisingly challenging  (Clark et al., 2019). (2) Timeframe Prediction, which requires predicting a specific timeframe for an event, and are essential for applications where knowing whether an event will happen or not without a timeframe is insufficient. (3) Quantity Estimation, which involves providing numerical estimates related to future events, e.g., economic indicators or trends.
@@ 3 Evaluating Predictions and Confidence
Evaluating Boolean Questions
For Boolean questions, where the answer space is $\mathcal{A}=\{\text{Yes},\text{No}\}$, prediction performance is evaluated using standard classification metrics, including accuracy and F1-score.
Confidence calibration is assessed using a modified version of the Brier score (Brier, 1950), which measures the mean squared error between predicted confidence and the gold confidence provided in the dataset, which we assume is provided, and represents the likelihood of an event occurring. The modified Brier score is defined as:
@@
where $C^{\text{pred}}_{i}$ is the model’s predicted confidence and $C^{\text{gold}}_{i}$ is the gold confidence. This modification ensures that models are evaluated based on their ability to match the likelihood of an event. Lower Brier scores indicate better calibration, reflecting how well the predicted confidence aligns with the likelihood of the event.
Evaluating Timeframe Prediction
@@
For confidence calibration, rely on the Continuous Ranked Probability Score (CRPS) (Matheson and Winkler, 1976), which is a generalisation of the Mean Absolute Error to probabilistic forecasts, and extend it to compare the predicted probability distribution with a gold distribution. Specifically, we assume that both the predicted and the gold confidence predictions follow Gaussian distributions, namely $\mathcal{N}(D^{\text{pred}}_{i},\sigma^{\text{pred}}_{i})$ and $\mathcal{N}(D^{\text{gold}}_{i},\sigma^{\text{gold}}_{i})$ respectively, where the standard deviations are computed as:
@@ 4.1 Data Source and Question Selection
FOReCAst is constructed from Metaculus,222www.metaculus.com. Examples in Appendix A. an online forecasting platform where forecasters submit probabilistic predictions to questions across various domains. Metaculus aggregates individual forecasts into a continuously updated community prediction, which is finalized just before resolution. Each question has predefined resolution criteria, ensuring verifiable outcomes. To ensure dataset reliability, we include only questions with a definitive resolution and at least 100 forecasts to maintain statistical reliability. Ambiguous or subjectively resolved questions are excluded, and we remove those whose outcomes depend on arbitrary or uncontrollable factors. These steps ensure that FOReCAst consists of high-quality, well-formed forecasting questions with verifiable outcomes. The dataset spans a diverse range of domains such as politics, economics, health, and technology, with questions distributed across a wide temporal range; detailed statistics are provided in Table 9.
@@ 4.2 Extracting Confidence from Crowdsourced Forecasts
Forecasting distinguishes between physical probabilities—objective likelihoods derived from statistical or scientific models—and human beliefs (Sanders, 1963) about future events. While physical probabilities can be useful, they are often unavailable, particularly for questions involving human behavior, economics, or sociopolitical outcomes. Instead, collective human forecasts offer a more practical confidence estimate, integrating expert reasoning, contextual knowledge, and evolving evidence. For instance, predicting a technological breakthrough depends more on expert assessment and current trends than on rigid probabilistic models. Therefore, confidence in FOReCAst is derived from Metaculus community forecasts, which aggregate predictions from a diverse pool of forecasters. While human predictions are sometimes incorrect, they still serve as a valuable proxy for uncertainty, as they reflect the best available reasoning given the information at the time.
Formal Definition of Gold Confidence
Gold confidence in FOReCAst is derived from the final Metaculus community prediction just before question resolution. Rather than directly using the predicted probability for the correct outcome, we compute a log score relative to a uniform baseline, following Metaculus’s established approach for evaluating probabilistic forecasts  [^3] https://www.metaculus.com/help/scores-faq/. This transformation ensures that confidence reflects not only correctness, but also the amount of informational gain over random guessing. It provides a consistent interpretation of confidence across tasks and scales, particularly when outcomes lie in continuous spaces.
For Boolean questions, where the human-forecasted probability for the correct outcome is $P^{\text{gold}}$, gold confidence is computed as:
@@
While this scoring formulation is tailored to Metaculus, it can be adapted or replaced if used with other forecasting platforms or scoring protocols. Our goal is to align confidence with meaningful deviations from uncertainty, while maintaining a consistent and interpretable signal across all tasks.
@@ 4.3 Dataset Statistics and Comparison
FOReCAst consists of 2256 forecasting questions, spanning domains such as politics, economics, science, and technology. Each question includes a resolved outcome, a gold confidence score, and a final Metaculus community forecast before resolution. To facilitate model development and evaluation, we split the dataset into 65% training, 10% validation, and 25% test. The full dataset statistics is shown in Appendix B.
@@ 5 Experiments on FOReCAst
5.1 Experimental Setup
Models.
We evaluate a diverse set of large language models (LLMs) with varying training data cutoffs, model sizes, and instruction tuning. To analyze the impact of knowledge recency, we group models by family and assume the cutoff date is the 1st of the stated month. The models used are shown in Table 3.
@@ 5.1 Experimental Setup
GPT-2 (Radford et al., 2019) GPT-2, GPT-2 XL 2017-12-01
Pythia (Biderman et al., 2023) 14M, 160M, 2.8B 2020-03-01
BLOOM (Scao et al., 2023) 560M, 7B1 2021-12-01
LLaMA (Touvron et al., 2023) LLaMA-7B 2022-08-01
OLMo (Groeneveld et al., 2024) 1B, 7B, 7B-Instruct 2023-03-01
OLMo-2 (OLMo et al., 2024) 7B, 7B-Instruct 2023-12-01
Inference.
We use 1-shot in-context learning to provide models with a structured example of how to answer forecasting questions. For instruction-tuned models, we add an extra line of instruction to align with their training paradigm. To ensure fair comparison, baseline prompts are kept minimal while maintaining clarity. Confidence is estimated using logit-based normalized heuristics. Full prompts and hyperparameters are provided in Appendix C and  Appendix D.
@@
Model N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
GPT2 401 0.58 0.37 0.42
GPT2-XL 401 0.67 0.32 0.45
Pythia-14m 343 0.59 0.14 0.62
Pythia-160m 343 0.61 0.26 0.53
Pythia-2.8b 343 0.55 0.39 0.44
Bloom-560m 263 0.49 0.41 0.45
Bloom-7b1 263 0.64 0.33 0.32
Llama-7b 226 0.57 0.34 0.52
OLMo-1B 188 0.23 0.16 0.75
OLMo-7B 188 0.21 0.17 0.82
OLMo-7B-Instruct 188 0.65 0.14 0.38
OLMo-2-7B 145 0.51 0.24 0.44
OLMo-2-7B-Instruct 145 0.59 0.30 0.41
Table 4: Performance of forecasting models on Boolean questions in FOReCAst. Reported metrics include the number of evaluated questions (N), accuracy, F1 score (both higher is better), and the Brier score (lower is better).
5.2 Results, Findings, and Discussions
Our experiments on the FOReCAst dataset reveal that forecasting remains a challenging task for current LLMs, particularly in estimating and expressing uncertainty. While models often achieve reasonable accuracy in their point predictions, their confidence calibration—measured by Brier score and CRPS—shows substantial variability. This indicates that uncertainty estimation should be treated as a distinct evaluation axis, not inferred from accuracy alone. Throughout the following sections, we offer interpretations and possible explanations for the observed behaviors.

### v4

@@ Abstract
Forecasting is an important task in many domains. However, existing forecasting benchmarks lack comprehensive confidence assessment, focusing on limited question types, and often consist of artificial questions that do not reflect real-world needs. To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), a benchmark that evaluates models’ ability to make predictions and their confidence in them. FOReCAst spans diverse forecasting scenarios involving Boolean questions, timeframe prediction, and quantity estimation, enabling a comprehensive evaluation of both prediction accuracy and confidence calibration for real-world applications. [^1] FOReCAst is publicly available at https://huggingface.co/datasets/MoyYuan/FOReCAst.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks. In parallel to these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023). Fact-oriented benchmarks, such as TruthfulQA (Lin et al., 2022), evaluate LLMs on factual correctness, focusing on tasks like retrieving and verifying facts that are already known.
@@
Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy. Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information. Several datasets have been introduced to evaluate LLMs’ forecasting capabilities. ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment. AutoCast (Zou et al., 2022) incorporates confidence intervals but not for its forecasting questions. Other datasets, such as ExpTime (Yuan et al., 2024a), are artificially generated from temporal knowledge graph, focusing on explainability.
@@
All of these aforementioned benchmarks overlook a crucial aspect of forecasting: confidence evaluation. Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive answers at the time they are asked. Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events. Moreover, miscalibrated confidence can lead to poor decision-making: overconfident yet incorrect forecasts may result in costly errors, while underconfident but accurate predictions can erode trust in the model. Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
@@
To address these issues we present FOReCAst: Future Outcome Reasoning and Confidence Assessment. Gold confidence scores are derived from aggregated human forecasts, reflecting community consensus over time. FOReCAst focuses on three distinct types of forecasting questions, as illustrated in Table 1: (1) Boolean Questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in 2025?" We conduct experiments using a range of models differing in size, training objectives, and training cutoff times, and explore multiple methods for estimating model confidence. Our results reveal that forecasting remains highly challenging for contemporary LLMs. Their confidence calibration is poor overll, and it has inconsistent correlation with predictive accuracy. While larger models sometimes improve performance, the effect is inconsistent and highly task-dependent.
Table 1: Forecasting questions with their resolutions and confidence scores.
@@
Boolean Question Will there be a Frontier AI lab in China before 2026? Yes 0.73
@@ 2 FOReCAst: Problem Formulation
In FOReCAst a system response consists of (1) a prediction that answers a question given the available information and (2) a confidence score in the prediction. This ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
Questions belong to three types. (1) Boolean Questions, which ask yes/no questions about the occurrence of future events (sometimes within a certain timeframe). Even if such questions are simple to evaluate, they can still be surprisingly challenging  (Clark et al., 2019). (2) Timeframe Prediction, which requires predicting a specific timeframe for an event, and are essential for applications where knowing whether an event will happen or not without a timeframe is insufficient. (3) Quantity Estimation, which involves providing numerical estimates related to future events.
@@ 3 Evaluating Predictions and Confidence
Boolean Questions
For Boolean questions, where the answer space is $\mathcal{A}=\{\text{Yes},\text{No}\}$, prediction performance is evaluated using standard classification metrics, including accuracy and F1-score. Confidence calibration is assessed using a modified version of the Brier score (Brier, 1950), which measures the mean squared error between predicted confidence and the gold confidence provided in the dataset, which we assume is provided, and represents the likelihood of an event occurring. The modified Brier score is defined as:
@@
where $C^{\text{pred}}_{i}$ is the model’s predicted confidence and $C^{\text{gold}}_{i}$ is the gold confidence. This modification ensures that models are evaluated based on their ability to match the likelihood of an event, with lower Brier scores indicating better calibration.
Timeframe Prediction
@@
For confidence calibration, we rely on the Continuous Ranked Probability Score (CRPS) (Matheson and Winkler, 1976), which is a generalisation of the Mean Absolute Error to probabilistic forecasts, and extend it to compare the predicted probability distribution with a gold distribution. Specifically, we assume that both the predicted and the gold confidence predictions follow Gaussian distributions $\mathcal{N}(D^{\text{pred}}_{i},\sigma^{\text{pred}}_{i})$ and $\mathcal{N}(D^{\text{gold}}_{i},\sigma^{\text{gold}}_{i})$ respectively, where the standard deviations are computed as:
@@ 4.1 Data Source and Question Selection
FOReCAst is constructed from Metaculus (www.metaculus.com, see examples in Appendix A), an online forecasting platform where users submit forecasts to questions across domains such as politics, economics, health, and technology. Metaculus aggregates these into a community prediction that is continuously updated until shortly before resolution. Each question has predefined resolution criteria, and the platform enforces guidelines to ensure that all predictions comply with them and are human-generated. We discard questions that Metaculus marks as ambiguous or those resolved in ways that make forecasting infeasible—for example, when the outcome falls outside the prediction range or when resolution criteria change after submission. Also, we only include questions with at least 100 forecasts to maintain statistical reliability.
@@ 4.2 Extracting Confidence from Crowdsourced Forecasts
In an ideal scenario, we would have both the resolutions and associated confidence scores available at the time a forecast question is posed. However, in practice, we only observe questions with their eventual resolutions and crowd-sourced human forecasts. To derive confidence scores we aggregate these human forecasts, which, despite occasional disagreements especially on non-trivial questions, provide informative signals about confidence. Human predictions can be sometimes incorrect, but they can still serve as a valuable proxy for uncertainty, as they reflect the best available reasoning given the information at the time. Metaculus ensures that these forecasts are manually submitted and fall within calibrated ranges, which encourages us to use them as proxies for evaluating model confidence. This is inline with other approaches to benchmark construction on tasks where humans disagree on labels, e.g. ChaosNLI (Nie et al., 2020; Plank, 2022).
Gold confidence in FOReCAst is derived from the final Metaculus community prediction just before question resolution. Following Metaculus’s established approach for evaluating probabilistic forecasts,  [^2] https://www.metaculus.com/help/scores-faq/ we compute a log score relative to a uniform baseline. This transformation ensures that confidence reflects not only correctness, but also the amount of informational gain over random guessing. It provides a consistent interpretation of confidence across tasks and scales, particularly when outcomes lie in continuous spaces. For Boolean questions, where the human-forecasted probability for the correct outcome is $P^{\text{gold}}$, gold confidence is computed as:
@@ 4.3 Dataset Statistics and Comparison
FOReCAst consists of 2256 forecasting questions, spanning domains such as politics, economics, science, and technology. Each question includes a resolved outcome, a gold confidence score, and a final Metaculus community forecast before resolution. The questions cover a broad temporal range, with resolutions spanning from March 2016 to the end of January 2025. To facilitate model development and evaluation, we split the dataset into 65% training, 10% validation, and 25% test. The full dataset statistics is shown in Appendix B. Note that we do not use the training split for model fine-tuning or prompt optimization; all models are evaluated in a one-shot setting.
@@
Extensibility and Long-Term Sustainability
FOReCAst is built with extensibility and sustainability in mind. Our data collection pipeline is fully automated, enabling continuous updates as new forecasting questions and outcomes become available. Although this release draws from Metaculus due to its high-quality resolution criteria and historical depth, the same pipeline can be adapted to other platforms such as Polymarket, Manifold Markets, or Good Judgment Open. By supporting multiple data sources and maintaining alignment with evolving real-world events, FOReCAst provides a sustainable foundation for benchmarking forecasting models over time. This extensibility not only ensures long-term utility, but also enables future work to study model generalization across domains, platforms, and temporal shifts.
@@ 5 Experiments on FOReCAst
5.1 Models
We evaluate a diverse set of large language models (LLMs) with varying training data cutoffs, model sizes, and instruction tuning. To analyze the impact of knowledge recency, we group models by family and use the stated month in Table 3 as the assumed cutoff. [^3] The exact cutoff date of Qwen 2.5 is unclear, with sources suggesting dates ranging from October 2023 to December 2023; we conservatively use 2023-12 as the assumed cutoff to avoid potential knowledge leakage.
@@ 5.1 Models
GPT-2 (Radford et al., 2019) GPT-2, GPT-2 XL 2017-12
Pythia (Biderman et al., 2023) 14M, 160M, 2.8B 2020-03
BLOOM (Scao et al., 2023) 560M, 7B1 2021-12
LLaMA (Touvron et al., 2023) LLaMA-7B 2022-08
OLMo (Groeneveld et al., 2024) 1B, 7B, 7B-Instruct 2023-03
OLMo-2 (OLMo et al., 2024) 7B, 7B-Instruct 2023-12
Qwen 2.5 (Yang et al., 2024) 1.5B-Instruct, 7B-Instruct 2023-12
LLaMa 3.1 (Grattafiori et al., 2024) 8B 2023-12
5.2 Inference
Each model is evaluated independently using 1-shot in-context learning. For every question, we sample $n=10$ outputs from the model using temperature-based decoding. Each output includes a prediction $v_{i}$ and an associated confidence score $c_{i}$, derived from the normalized logits. Specifically, we compute the log-probability of each generated token using $\log p_{i,j}=\log\text{softmax}(z_{i,j})[x_{i,j}]$, where $z_{i,j}$ is the model’s pre-softmax logit vector at position $j$ for sample $i$, and $x_{i,j}$ is the generated token index. The confidence score $c_{i}$ is then defined as the average log-probability across all tokens: $c_{i}=\frac{1}{T}\sum_{j=1}^{T}\log p_{i,j}$, where $T$ is the length of the generated sequence. We observe low variance across the $n$ samples, with most models producing consistent predictions and confidence values.
@@ 5.2 Inference
Prompts are minimal and structured to ensure consistency across models. Each input includes a forecasting question and a single demonstration example. Instruction-tuned models receive an additional natural-language instruction to align with their training paradigm. Full prompt templates and decoding hyperparameters are provided in Appendix C and Appendix D.
5.3 Results
Our experiments on the FOReCAst dataset reveal that forecasting remains a challenging task for current LLMs, particularly in estimating and expressing uncertainty. While models often achieve reasonable accuracy in their point predictions, their confidence calibration, measured by Brier score and CRPS, shows substantial variability. This indicates that confidence estimation should be treated as a distinct evaluation axis, not inferred from accuracy alone. Table 4 reports results on a shared question set to ensure fair comparison across models, while Table 5 evaluates each model on all questions available up to its respective training cutoff as a view of each model’s forecasting potential.
Table 4: Combined forecasting performance for cutoff date 2023-12, ensuring same questions for all models. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.59 0.31 0.37 1.00 1.00 0.25 0.73 0.72
GPT2-XL 0.71 0.30 0.42 1.00 1.00 0.04 0.69 0.64
Pythia-14m 0.62 0.14 0.65 1.00 1.00 0.25 0.85 0.83
Pythia-160m 0.67 0.31 0.51 1.00 1.00 0.02 0.67 0.65
Pythia-2.8b 0.52 0.29 0.45 1.00 1.00 0.12 0.73 0.69
Bloom-560m 0.52 0.38 0.41 1.00 1.00 0.03 0.67 0.65
Bloom-7b1 0.68 0.36 0.30 1.00 1.00 0.11 0.78 0.75
Llama-7b 0.54 0.27 0.55 0.96 0.94 0.06 0.62 0.59
OLMo-1B 0.23 0.21 0.75 1.00 1.00 0.22 0.76 0.74
OLMo-7B 0.22 0.23 0.83 1.00 1.00 0.22 0.82 0.81
OLMo-7B-Inst 0.68 0.16 0.36 1.00 1.00 0.27 0.81 0.78
OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
OLMo-2-7B-Inst 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
Qwen-2.5-1.5B-Inst 0.67 0.20 0.32 1.00 1.00 0.14 0.74 0.72
Qwen-2.5-7B-Inst 0.68 0.08 0.25 1.00 1.00 0.07 0.55 0.53
Llama-3.1-8B 0.08 0.00 0.99 1.00 1.00 0.02 0.58 0.54
Table 5: Forecasting performance across all tasks in FOReCAst, evaluated using questions available to each model based on its training cutoff date. As a result, the number of evaluated questions ($N$) may vary across models. Boolean question metrics include Accuracy, F1, and Brier score (higher is better for Accuracy/F1, lower is better for Brier). Quantity Estimation and Timeframe Prediction use lower-is-better metrics. APE: Absolute Percentage Error, MAE: Mean Absolute Error, CRPS: Continuous Ranked Probability Score, ADE: Absolute Days Error.
Model Boolean Questions Quantity Estimation Timeframe Prediction
N Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$) N ADE ($\downarrow$) CRPS ($\downarrow$)
GPT2 401 0.58 0.37 0.42 81 0.23 0.87 0.86 26 0.99 0.99
GPT2-XL 401 0.67 0.32 0.45 81 0.03 0.75 0.72 26 1.00 1.00
Pythia-14m 343 0.59 0.14 0.62 77 0.17 0.88 0.85 25 1.00 1.00
Pythia-160m 343 0.61 0.26 0.53 77 0.03 0.78 0.76 25 1.00 1.00
Pythia-2.8b 343 0.55 0.39 0.44 77 0.07 0.81 0.79 25 0.97 0.96
Bloom-560m 263 0.49 0.41 0.45 43 0.05 0.73 0.71 12 1.00 1.00
Bloom-7b1 263 0.64 0.33 0.32 43 0.06 0.75 0.72 12 1.00 1.00
Llama-7b 226 0.57 0.34 0.52 32 0.06 0.64 0.61 10 0.98 0.98
OLMo-1B 188 0.23 0.16 0.75 23 0.21 0.79 0.76 6 1.00 1.00
OLMo-7B 188 0.21 0.17 0.82 23 0.22 0.84 0.82 6 1.00 1.00
OLMo-7B-Inst 188 0.65 0.14 0.38 23 0.24 0.80 0.77 6 1.00 1.00
OLMo-2-7B 145 0.51 0.24 0.44 20 0.02 0.65 0.62 4 1.00 1.00
OLMo-2-7B-Inst 145 0.59 0.30 0.41 20 0.09 0.60 0.57 4 0.87 0.82
Qwen-1.5B-Inst 145 0.67 0.20 0.32 20 0.14 0.74 0.72 4 1.00 1.00
Qwen-7B-Inst 145 0.68 0.08 0.25 20 0.07 0.55 0.53 4 1.00 1.00
Llama-3.1-8B 145 0.08 0.00 0.99 20 0.02 0.58 0.54 4 1.00 1.00